# Explicit bounds for the graphicality of the prime gap sequence

### v1

@@ 1.1. DPG-process
The pair $(G_{n},d_{n+1})$ is called DPG-graphic if $G_{n}$ contains a matching of size $d_{n+1}/2$. By [8, Theorem 2.4], the sequence $D_{n}\circ d_{n+1}$ is graphic if and only if there exists a realization $G_{n}$ of $D_{n}$ such that the pair $(G_{n},d_{n+1})$ is DPG-graphic. In order to construct an infinite DPG process, one needs the requirement that every realization of $D_{n}$ contains a matching of size $d_{n+1}/2$. While this holds trivially for constant sequences (e.g., growing regular graphs), general sequences require careful analysis of matchings in $G_{n}$. To achieve this, we will rely on Vizing’s theorem and its consequence stated in Lemma 2.8.
@@ 2.1. Some explicit estimates
Let $H_{0}\in\mathbb{R}$ be such that the RH is known to be true up to $H_{0}$. That is, all zeroes $\beta+\mathrm{i}\gamma$ of the Riemann zeta-function with $|\gamma|<H_{0}$ have $\beta=1/2$. Moreover, all of these zeroes are simple. Currently, we may take $H_{0}=3\times 10^{12}$ (see Platt and Trudgian[13, Theorem 1]). For $3/5\leq\sigma\leq 1$ and $T>H_{0}$, we have the following zero-density estimate (see Kadiri, Lumley and Ng [11, Section 5, Table 1])
@@
By the zero-free region of Ford [9, Theorem 5], we have that for $|t|>3$, every non-trivial zero $\beta+\mathrm{i}t$ of the Riemann zeta function satisfies
@@
The case $t>50$ follows from Dudek [4, Lemma 2.6]. For $10<t\leq 50$, the inequality holds by direct verification of zeros of $\zeta(s)$ tabulated in the LMFDB database (https://www.lmfdb.org/zeros/zeta/). ∎
@@
Lemma 2.4 (Dudek [4, Lemma 2.8]).
@@
This proof follows the framework laid out in Section 6 of [7] with some necessary modifications. We start with the truncated Perron’s formula, as presented in [7, Eq. (19)]. To avoid the potential issue that $T$ might lie close to the ordinate of a zero, we choose some $T_{0}\in(T-1,T+1)$ according to Lemma 2.4. For a non-integer $x>2$ and $c=1+1/\log x$, we have
@@
For $S_{1}$, we use the second term in the $\min$ function. By [7, Eq. (23)], we have
@@
Finally for $S_{4}$, there are two $n$’s satisfying |xn|<1|x-n|<1. We use the first term in the $\min$ function, and by [7, Eq. (20)], the contribution is
@@ 3. First moment of large prime gaps
Lemma 3.1 is also established in [7, Section 5]. Here, we present a reformulation and a minor modification to it. Specifically, the conclusion of Lemma 2.1 ensures validity for $T<x$, which allows us to take a smaller imaginary part.
The main task now is to give the non-trivial estimate of Integral $\mathcal{I}$ explicitly. This can be done by following the same idea as in Saffari and Vaughan [15, Lemma 5].
@@ 4.2. Case 2: $k_{N}\leq\sqrt{p_{n}}$
Using the fact that there is at least one prime between $n^{140}$ and $(n+1)^{140}$ for all $n\geq 1$ (see Cully-Hugill and Johnston [2, Theorem 1.3]), we crudely bound
$M(p_{n})\leq 280\,p_{n}^{139/140}\left(1+\frac{2}{p_{n}^{1/140}}\right)^{139}\leq\frac{p_{n}}{3\log^{2}p_{n}}$}
@@ 5. About the DPG-process
We have proved that the sequence of first $n$ prime gaps (with $p_{0}=1$) is graphic for $n\geq\exp\exp(30.49)$. Let $G_{n}$ be the realization. By Erdős et al.[8, Theorem 2.5], to prove that $(p_{\ell}-p_{\ell-1})_{\ell=1}^{n}$ is DPG-graphic, it suffices to show that $G_{n}$ has $(p_{n+1}-p_{n})/2$ independent edges. Applying Lemma 2.8, we are required to exhibit an integer $N\geq 2$ satisfying
@@
for all $n\geq n_{0}$ (This is [7, Eq. (11)]). In [4], the author shows that for $x>\exp\exp(33.3)$ one has the bound $p_{n+1}-p_{n}<3x^{2/3}+3x^{1/3}+2$ when $x<p_{n}<2x$. Choose $\alpha=1/13$ such that
@@
Acknowledgements The authors would like to thank Prof. Gergely Harcos for suggesting the problem and his valuable suggestions on this topic. K. Aggarwal thanks the Rényi Institute for funding and support where most of the work was done. R. Frot also thanks the Rényi and the Erdős Center for funding and support. H. Gou and H. Wang are also grateful to Prof. Jianya Liu for his help and encouragement. H. Wang also gratefully acknowledges the Rényi Institute for its excellent research environment, which enabled the preparation of a portion of this work, as well as the China Scholarship Council for supporting her studies there.

### v2

@@ 1.1. DPG-process
The pair $(G_{n},d_{n+1})$ is called DPG-graphic if $G_{n}$ contains a matching of size $d_{n+1}/2$. By [9, Theorem 2.4], the sequence $D_{n}\circ d_{n+1}$ is graphic if and only if there exists a realization $G_{n}$ of $D_{n}$ such that the pair $(G_{n},d_{n+1})$ is DPG-graphic. In order to construct an infinite DPG process, one needs the requirement that every realization of $D_{n}$ contains a matching of size $d_{n+1}/2$. While this holds trivially for constant sequences (e.g., growing regular graphs), general sequences require careful analysis of matchings in $G_{n}$. To achieve this, we will rely on Vizing’s theorem and its consequence stated in Lemma 2.8.
@@ 2.1. Some explicit estimates
Let $H_{0}\in\mathbb{R}$ be such that the RH is known to be true up to $H_{0}$. That is, all zeroes $\beta+\mathrm{i}\gamma$ of the Riemann zeta-function with $|\gamma|<H_{0}$ have $\beta=1/2$. Moreover, all of these zeroes are simple. Currently, we may take $H_{0}=3\times 10^{12}$ (see Platt and Trudgian[15, Theorem 1]). For $3/5\leq\sigma\leq 1$ and $T>H_{0}$, we have the following zero-density estimate (see Kadiri, Lumley and Ng [13, Section 5, Table 1])
@@
By the zero-free region of Ford [10, Theorem 5], we have that for $|t|>3$, every non-trivial zero $\beta+\mathrm{i}t$ of the Riemann zeta function satisfies
@@
The case $t>50$ follows from Dudek [5, Lemma 2.6]. For $10<t\leq 50$, the inequality holds by direct verification of zeros of $\zeta(s)$ tabulated in the LMFDB database (https://www.lmfdb.org/zeros/zeta/). ∎
@@
Lemma 2.4 (Dudek [5, Lemma 2.8]).
@@
This proof follows the framework laid out in Section 6 of [8] with some necessary modifications. We start with the truncated Perron’s formula, as presented in [8, Eq. (19)]. To avoid the potential issue that $T$ might lie close to the ordinate of a zero, we choose some $T_{0}\in(T-1,T+1)$ according to Lemma 2.4. For a non-integer $x>2$ and $c=1+1/\log x$, we have
@@
For $S_{1}$, we use the second term in the $\min$ function. By [8, Eq. (23)], we have
@@
Finally for $S_{4}$, there are two $n$’s satisfying |xn|<1|x-n|<1. We use the first term in the $\min$ function, and by [8, Eq. (20)], the contribution is
@@
Remark 2.
Note that Lemma 2.5 gives an explicit version of a result due to Goldston [11, Eq. (4)]; see also another explicit version in Cully-Hugill [2, Theorem 2] for half odd integers.
@@ 3. First moment of large prime gaps
Lemma 3.1 is also established in [8, Section 5]. Here, we present a reformulation and a minor modification to it. Specifically, the conclusion of Lemma 2.1 ensures validity for $T<x$, which allows us to take a smaller imaginary part.
The main task now is to give the non-trivial estimate of Integral $\mathcal{I}$ explicitly. This can be done by following the same idea as in Saffari and Vaughan [17, Lemma 5].
@@ 4.2. Case 2: $k_{N}\leq\sqrt{p_{n}}$
Using the fact that there is at least one prime between $n^{90}$ and $(n+1)^{90}$ for all $n\geq 1$ (see Cully-Hugill and Johnston [3, Theorem 1.4]), we crudely bound
$M(p_{n})\leq 180\,\left(p_{n}^{1/90}+2\right)^{89}\leq\frac{p_{n}}{3\log^{2}p_{n}}$}
@@ 5. About the DPG-process
We have proved that the sequence of first $n$ prime gaps (with $p_{0}=1$) is graphic for $n\geq\exp\exp(30.49)$. Let $G_{n}$ be the realization. By Erdős et al.[9, Theorem 2.5], to prove that $(p_{\ell}-p_{\ell-1})_{\ell=1}^{n}$ is DPG-graphic, it suffices to show that $G_{n}$ has $(p_{n+1}-p_{n})/2$ independent edges. Applying Lemma 2.8, we are required to exhibit an integer $N\geq 2$ satisfying
@@
for all $n\geq n_{0}$ (This is [8, Eq. (11)]). In [5], the author shows that for $x>\exp\exp(33.3)$ one has the bound $p_{n+1}-p_{n}<3x^{2/3}+3x^{1/3}+2$ when $x<p_{n}<2x$. Choose $\alpha=1/13$ such that
@@
Acknowledgements. The authors would like to thank Prof. Gergely Harcos for suggesting the problem and his valuable suggestions on this topic. We also thank Daniel Johnston for pointing out the references [2, 3]. K. Aggarwal thanks the Rényi Institute for funding and support where most of the work was done. R. Frot also thanks the Rényi and the Erdős Center for funding and support, which were provided in part by the Rényi Intézet Lendület Automorphic Research Group. H. Gou and H. Wang are grateful to Prof. Jianya Liu for his help and encouragement. H. Wang also gratefully acknowledges the Rényi Institute for its excellent research environment, which enabled the preparation of a portion of this work, as well as the China Scholarship Council for supporting her studies there.